OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

53 :D

OpenStudy (anonymous):

Okay so you can find the answer two ways. 1. 2*6=12 65-12=53 2.-2*6=-12+65=53